OpenStudy (angelwings996):

Verify the Pythagorean Identity. 1 + cos^2(theta) = csc^2(theta)

OpenStudy (angelwings996):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

you sure it's cos ? and not cot ?

OpenStudy (angelwings996):

Yeah its cot, sorry about that

jimthompson5910 (jim_thompson5910):

that's ok

jimthompson5910 (jim_thompson5910):

use the fact that cot(x) = cos(x)/sin(x)

OpenStudy (angelwings996):

Okay

jimthompson5910 (jim_thompson5910):

doing so will give you \[\large 1 + \cot^2(\theta) = \csc^2(\theta)\] \[\large 1 + \frac{\cos^2(\theta)}{\sin^2(\theta)} = \csc^2(\theta)\] what's your next step?

OpenStudy (angelwings996):

Uhmm... would you do something with the 1 ? Im not completly sure

jimthompson5910 (jim_thompson5910):

try combine the 1 with \(\large \frac{\cos^2(\theta)}{\sin^2(\theta)} \)

jimthompson5910 (jim_thompson5910):

so you need to rewrite 1 as a fraction with the denominator \(\large \sin^2(\theta)\) since that's the LCD

jimthompson5910 (jim_thompson5910):

then you can combine the fractions

OpenStudy (angelwings996):

cot^2(x) + 1 ?

jimthompson5910 (jim_thompson5910):

idk how you got cot

OpenStudy (angelwings996):

Hold on, Im thinking

jimthompson5910 (jim_thompson5910):

alright

OpenStudy (angelwings996):

Okay so I need to have 1 with sin in the denominator?

jimthompson5910 (jim_thompson5910):

sin^2 actually

jimthompson5910 (jim_thompson5910):

since that's the denominator of \(\large \frac{\cos^2(\theta)}{\sin^2(\theta)} \)

OpenStudy (angelwings996):

\[\frac{ \cos ^{2}\Theta }{ \sin ^{2}1\Theta } ?\]

OpenStudy (angelwings996):

or would there be another 1 somewhere

jimthompson5910 (jim_thompson5910):

if the denominator was x, then we can write 1 as x/x

OpenStudy (angelwings996):

Oh, then just put 1 as a fraction with sin^2 ?

jimthompson5910 (jim_thompson5910):

yes, what do you get

OpenStudy (angelwings996):

\[\frac{ \cos ^{2}\theta }{ \frac{ \sin ^{2} }{ 1 } \theta} ? \]

jimthompson5910 (jim_thompson5910):

no

jimthompson5910 (jim_thompson5910):

\[\large 1 = \frac{\sin^2(\theta)}{\sin^2(\theta)}\]

jimthompson5910 (jim_thompson5910):

ie...anything over itself is always 1

OpenStudy (angelwings996):

Ohh

jimthompson5910 (jim_thompson5910):

so what this means is that \[\large 1 + \frac{\cos^2(\theta)}{\sin^2(\theta)} = \csc^2(\theta)\] turns into \[\large \frac{\sin^2(\theta)}{\sin^2(\theta)} + \frac{\cos^2(\theta)}{\sin^2(\theta)} = \csc^2(\theta)\]

OpenStudy (angelwings996):

Okay I follow you

jimthompson5910 (jim_thompson5910):

so what's next

OpenStudy (angelwings996):

I add? Or do I need to cancel out the sin^2 (theta) ?

jimthompson5910 (jim_thompson5910):

you add the fractions

OpenStudy (angelwings996):

would it be sin^2+cos^2x/sin^2x = csc^2x ?

jimthompson5910 (jim_thompson5910):

keep going

OpenStudy (angelwings996):

Or does the top part need to be sin^2x + cos^2x ?

jimthompson5910 (jim_thompson5910):

it should be \[\large \frac{\sin^2(\theta)+\cos^2(\theta)}{\sin^2(\theta)} = \csc^2(\theta)\]

jimthompson5910 (jim_thompson5910):

because remember that a/c + b/c = (a+b)/c

OpenStudy (angelwings996):

Okay, now do I cancel out sin^2x ?

jimthompson5910 (jim_thompson5910):

no

jimthompson5910 (jim_thompson5910):

what does the top simplify to?

jimthompson5910 (jim_thompson5910):

\[\large \sin^2(\theta)+\cos^2(\theta) = ??\]

OpenStudy (angelwings996):

Uh.. it equals 1 ?

OpenStudy (angelwings996):

Yeah it equals 1 right?

jimthompson5910 (jim_thompson5910):

yep

jimthompson5910 (jim_thompson5910):

\[\large \sin^2(\theta)+\cos^2(\theta) = 1\]

jimthompson5910 (jim_thompson5910):

you now have \[\large \frac{1}{\sin^2(\theta)} = \csc^2(\theta)\]

OpenStudy (angelwings996):

Okay

jimthompson5910 (jim_thompson5910):

do you see what's next?

OpenStudy (angelwings996):

No

jimthompson5910 (jim_thompson5910):

what is csc

OpenStudy (angelwings996):

Idk...

jimthompson5910 (jim_thompson5910):

csc(x) = 1/sin(x) by definition of cosecant

jimthompson5910 (jim_thompson5910):

so we have 1/(sin^2(x)) which will turn into what?

OpenStudy (angelwings996):

csc^2x ?

jimthompson5910 (jim_thompson5910):

good

jimthompson5910 (jim_thompson5910):

\[\large \frac{1}{\sin^2(\theta)} = \csc^2(\theta)\] turns into \[\large \csc^2(\theta) = \csc^2(\theta)\]

jimthompson5910 (jim_thompson5910):

the two sides of the last equation are identical, so we have confirmed the identity

jimthompson5910 (jim_thompson5910):

Here are all the steps in one post \[\large 1 + \cot^2(\theta) = \csc^2(\theta)\] \[\large 1 + \frac{\cos^2(\theta)}{\sin^2(\theta)} = \csc^2(\theta)\] \[\large \frac{\sin^2(\theta)}{\sin^2(\theta)} + \frac{\cos^2(\theta)}{\sin^2(\theta)} = \csc^2(\theta)\] \[\large \frac{\sin^2(\theta)+\cos^2(\theta)}{\sin^2(\theta)} = \csc^2(\theta)\] \[\large \frac{1}{\sin^2(\theta)} = \csc^2(\theta)\] \[\large \csc^2(\theta) = \csc^2(\theta)\] the two sides of the last equation are identical, so we have confirmed the identity

OpenStudy (angelwings996):

Okay thank you so much!

jimthompson5910 (jim_thompson5910):

you're welcome